3.6 (10 pts) Solve the differential equation y = (4x +y – 5)2 by using appropriate transforma- tion.
Answer #1

3) Given that differenti al qualion y =(x + y - 52.0 Let 4x47 5= v be the transformation. Differentiating wertx on both sides22 (AX+Y-5) = tan (21 +2C) Ź 4x+y 5 = 2 tan (2x+20) Ź y 600 = 5- 4% +2tan (2X+20), This is the required solution of the given